2) For circuit A5-1 below, given that the input voltage to transformer AB is 12.4 KVZ0°, and Z is 300215° solve for I Ie Solve for the apparent power S delivered to the load Z. Solve for the the power factor of the load. Solve for the apparent power S at the primary side of trans former AB. What is the efficiency of the system in delivering real power to the load? (7 marks) 3) For circuit A5-1 below, given that the load is operating at 3 MVA, with a power factor of 0.9 lagging, and the voltage supplied to the load is 33.9KVZ0°, solve for I, Ia, Ic. Solve for the power delivered at the primary side of transformer AB, and power factor at the source. What is the efficiency of the system in delivering real power to the load? (6 marks) (Note: Circuit is single phase.) A5-1 XBc RBC R Xac IC z X AB2 XAB RA RAB IB
Transformer AB: 12.4/225kV Transformer BC: 225/34.6kV RвCI 3 50 RABI 6 ХАвт 3 4 RAB2 2000 Хвся %3 10 Rвcz 3D 0.1250 ХАВ2 150 Хвса 3 0.250
Answer #1

First the entire circuit is drawn by transferring all impedances to primary of transformer AB. To transfer impedances, multiply them by square of voltage ratio. To transfer current, multiply by inverse of voltage ratio.

RABI= 60 XAE 42 A82 XA82= 15X(12-4 225 R- 4 2 62 2a5 O.06 2 O.046 2 aa5 225 1O.015 SO.003 22 34.6 Rac2-0125x 22X 252 XBC2= O-
6- 091tj4-081)R ueeer I V A ep 1a 4X103 Gtoery. (6.09/ +j4.031)+38-531Z15. a73 3352-17.972° A- elf18 4p . TA= 2구2-3352-17.978
Load M VA=3Z csa.9 =3 L25.842 325.84210 33.9 X103 88.5-25.84 A A82.52-25842 A 0.P7-41 P A c BX4 3175-25 842 BX225 A Puray cot
383.051 ceo25-842 383 aslceo2S-842+6-071 X1001. -98.264